OpenStudy (anonymous):

WILL FAN AND MEDAL!!!! Find f(5) for this sequence: f(1) = 2 and f(2) = 5, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

jimthompson5910 (jim_thompson5910):

do you know what the notation " f(n) = f(1) + f(2) + f(n - 1)" means ?

OpenStudy (anonymous):

no sadly... :( ive been studying on the lesson for this all day and i have yet to figure it out

jimthompson5910 (jim_thompson5910):

we have f(1) = 2 and f(2) = 5 and that's given to us

jimthompson5910 (jim_thompson5910):

now let's say we want to find f(3)

jimthompson5910 (jim_thompson5910):

for f(3), that's basically f(n) when n = 3 agreed?

OpenStudy (anonymous):

ok hold on a sec i gots to read it first.... lol

OpenStudy (anonymous):

ok yea agreed

jimthompson5910 (jim_thompson5910):

so what we do is replace every copy of n with 3

jimthompson5910 (jim_thompson5910):

f(n) = f(1) + f(2) + f(n - 1) f(3) = f(1) + f(2) + f(3 - 1) f(3) = f(1) + f(2) + f(2)

jimthompson5910 (jim_thompson5910):

what's next?

OpenStudy (anonymous):

hold on i had to change my shirt cuz im gettin cold so yea...

OpenStudy (anonymous):

um... you add the f(1)+f(2)+f(2)?

jimthompson5910 (jim_thompson5910):

before you add, why not replace f(1) and f(2) with what they are equal to?

jimthompson5910 (jim_thompson5910):

f(1) = 2 f(2) = 5

OpenStudy (anonymous):

m my question is that how are u figuring out how to get what they are equal to...

jimthompson5910 (jim_thompson5910):

that information of f(1) = 2 and f(2) = 5 is given at the very top

jimthompson5910 (jim_thompson5910):

so whoever wrote this problem just made it up and gave it to you

OpenStudy (anonymous):

ooohhhh i see now ok so yea just a sec imma figure it out...

OpenStudy (anonymous):

ok so f(3)=12?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

f(1) = 2 f(2) = 5 f(3) = 12

jimthompson5910 (jim_thompson5910):

how do we find f(4) ?

OpenStudy (anonymous):

idk im stuck....

OpenStudy (jdoe0001):

hint: \(\large { f_1=2\qquad f_2=5\qquad f_{\color{brown}{ n}}=f_1+f_2+f_{\color{brown}{ n-1}}\qquad {\color{brown}{ n}}>2\\ \begin{array}{ccllll} term\quad f(n)&value \\\hline\\ 1&2\\ 2&5\\ {\color{brown}{ 3}}&f_1+f_2+f_{\color{brown}{ 3-1}}\\ &f_1+f_2+f_2\\ &2+5+5\\ 4&f_1+f_2+f_{\color{brown}{ n-1}}\\ 5&f_1+f_2+f_{\color{brown}{ n-1}} \end{array} }\)

jimthompson5910 (jim_thompson5910):

replace every copy of n in f(n) = f(1) + f(2) + f(n - 1) with 4

OpenStudy (jdoe0001):

ditto

jimthompson5910 (jim_thompson5910):

and see what happens

OpenStudy (jdoe0001):

\(\large { \begin{array}{ccllll} term\quad f(n)&value \\\hline\\ 1&2\\ 2&5\\ {\color{brown}{ 3}}&f_1+f_2+f_{\color{brown}{ 3-1}}\\ &f_1+f_2+f_2\\ &2+5+5\\ {\color{brown}{ 4}}&f_1+f_2+f_{\color{brown}{ 4-1}}\\ {\color{brown}{ 5}}&f_1+f_2+f_{\color{brown}{ 5-1}} \end{array} }\)

OpenStudy (anonymous):

ok hold on.... does the f2 and all in the work is f(2)?

jimthompson5910 (jim_thompson5910):

yeah \(\large f_{2}\) and f(2) are effectively the same thing just in different notation

OpenStudy (anonymous):

ok just makin sure

jimthompson5910 (jim_thompson5910):

f(2) is easier to write in plain text, so we'll stick with that

OpenStudy (anonymous):

ok hold up a sec

OpenStudy (anonymous):

f(4) = f(1) + f(2) + f(4 - 1)? and then f(4) = 2 + 5 + 3 ?

jimthompson5910 (jim_thompson5910):

f(4) = f(1) + f(2) + f(4 - 1) f(4) = f(1) + f(2) + f(3) f(4) = ???

jimthompson5910 (jim_thompson5910):

notice how I wrote f(3) and not just 3 all by itself

jimthompson5910 (jim_thompson5910):

so you can see how f(4) depends on f(3). You need to know the previous value of f(3) to get the value of f(4)

OpenStudy (anonymous):

oohhhh ok yea i see now but i cant believe i have to do all this work for one freakin problem.... i miss when we had 5*10=whatever but getting off track yea i see how they depend on each other

jimthompson5910 (jim_thompson5910):

it's not too bad once you get to know the problem better

OpenStudy (anonymous):

oh and i have one question... do u do flvs and if so are u in 9th or 10th grade?

jimthompson5910 (jim_thompson5910):

no I'm not in flvs

jimthompson5910 (jim_thompson5910):

what are you getting for f(4) ?

OpenStudy (anonymous):

f(4)=19?....

jimthompson5910 (jim_thompson5910):

f(4) = f(1) + f(2) + f(4 - 1) f(4) = f(1) + f(2) + f(3) f(4) = 2 + 5 + 12 f(4) = 19 looks good

jimthompson5910 (jim_thompson5910):

f(1) = 2 f(2) = 5 f(3) = 12 f(4) = 19

OpenStudy (anonymous):

ok my other question is how in the world did u get two medals already when only one other person was in the problem for a sec? and yea

jimthompson5910 (jim_thompson5910):

someone else gave it to me I think

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

well i am goin to medal u right now cuz u are great at helpin people and there needs to be more people in life like u! :)

jimthompson5910 (jim_thompson5910):

thanks

jimthompson5910 (jim_thompson5910):

hopefully you see how to get f(5) ?

OpenStudy (anonymous):

i do but just to make sure i need to do this f(n) = f(1) + f(2) + f(3) + f(4) + f(n-1) and replace the n with 5?

jimthompson5910 (jim_thompson5910):

you always go back to the original equation they give you, which is f(n) = f(1) + f(2) + f(n - 1)

jimthompson5910 (jim_thompson5910):

plug n = 5 into there

OpenStudy (anonymous):

ok so f(5) = f(1) + f(2) + f(3) + f(4) + f(5 - 1)?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

f(n) = f(1) + f(2) + f(n - 1) f(5) = f(1) + f(2) + f(5 - 1)

jimthompson5910 (jim_thompson5910):

I'm going back to the original equation they gave you

OpenStudy (anonymous):

oh so i need to use the original equation to find all the f(1 through 5)?

jimthompson5910 (jim_thompson5910):

well except for f(1) and f(2) because those are given to you

jimthompson5910 (jim_thompson5910):

but yeah

OpenStudy (anonymous):

ok yea so f(5)= f(1) + f(2) + f(5 - 1) ?

jimthompson5910 (jim_thompson5910):

yes and you will keep going

OpenStudy (anonymous):

well 5 is the number i need to find so yea

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

ok hold on a sec

OpenStudy (anonymous):

f(5)= f(1) + f(2) + f(5 - 1) f(5)= f(1) + f(2) + f(4) f(5)= 2 + 5 + 19 f(5)= 26?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

:D !!

OpenStudy (anonymous):

Thanks so much!!! imma fan u too cuz ur amazing thanks so much for helping me... u know... its hard to come by people like u.. thanks so much!! :D :) 11

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

thats suppose to be !! at the end lol

OpenStudy (anonymous):

but yea thanks so much and i guess this is where we split paths so i will see u later my friend!

jimthompson5910 (jim_thompson5910):

have a good day

OpenStudy (anonymous):

u too! bye!